Find zeros of f(z)=sin hz


1
Expert's answer
2021-09-07T18:46:58-0400

Zeros of function "f(z)=sin\,hz" can be found from equation "sin\,hz=0". They are: "z=\frac{2\pi n}{h}", "n\in{\mathbb{Z}}".